jhonyy9 (jhonyy9):

- resulted from prims subject : - a right triangle can be formed by length sides of prims ? what mean that need being true the pythagora's c^2 = a^2 +b^2 for a,b and c prim numbers - any idea,opinion ?

jhonyy9 (jhonyy9):

@ganeshie8

jhonyy9 (jhonyy9):

@Kainui

jhonyy9 (jhonyy9):

so than not - any proof of this why ?

OpenStudy (welshfella):

all 3 numbers would have to be odd. Is this possible?

OpenStudy (welshfella):

Oh - I suppose one of them could be 2.....

jhonyy9 (jhonyy9):

why all 3 - one from these can being 2 - so what is even - or not ?

OpenStudy (welshfella):

yea i agree. Its a tricky one , this.

OpenStudy (welshfella):

I would think its highly unlikely one is 2 . If you see a list of these triples none of the early ones have a 2 in them and when we get to the higher numbers a 2 is not possible

OpenStudy (welshfella):

if a and b are both odd we have (2n + 1^2 + (2n + 1)^2 = 8n^2 + 8n + 2 which must be even

OpenStudy (welshfella):

is that valid?

OpenStudy (welshfella):

- oh - only if both numbers are the same , of course! lol

OpenStudy (welshfella):

however (2n + 1)^2 + (3n + 1)^2 would also be even...

jhonyy9 (jhonyy9):

on this website i have got the right explination and the proof of this my question - https://www.mathsisfun.com/numbers/pythagorean-triples.html

ganeshie8 (ganeshie8):

\[c^2 = a^2 +b^2 \implies a^2 = (c+b)(c-b)\] Since \(a\) is prime, we get two cases : Case 1) \(c+b = c-b = a\) Case 2) \(c-b=1\) and \(c+b=a^2\) Both happen to be impossible.

jhonyy9 (jhonyy9):

@ganeshie8 yes this is nice very and easy understanding - i agree it

jhonyy9 (jhonyy9):

so and thank you

jhonyy9 (jhonyy9):

@phi your opinion please - ? thank you in advance

OpenStudy (kainui):

Awesome! This reminds me of a particularly fun problem you might enjoy @jhonyy9 What are all the primes p, q, and r that satisfy this equation: \[p+q^2=r^2\]

jhonyy9 (jhonyy9):

sorry but i think you missed the exponent of p - or not ?